OpenStudy (anonymous):

Select the system of equations that corresponds to the given graph. a graph of two lines. The first passes through the points negative 1 comma 0 and 0 comma negative 2 and the second passes through the points negative 2 comma 0 and 0 comma negative 4. -2x - y = 4 2x + y = -2 -8x + 4y = -16 -4x + 2y = -8 x - 2y = 6 -2x + y = 4 2x + 3y = 6 3x + 2y = 4

OpenStudy (goformit100):

x - 2y = 6 -2x + y = 4